# Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need

### v2

@@ Appendix A.3 Senators of the 117th US Congress
Appendix Appendix B Constructing and submitting prompts to LLMs
@@ Appendix B.1 OpenAI API
  • GPT-4 Turbo (gpt-4-0125-preview) and GPT 3.5 Turbo (gpt-3.5-turbo-0125): We used the following command (a JSON mode is available for these models but not for the earlier models):
@@ Appendix B.2 Mistral AI API
  • Models published in 2023: open-mistral-7b,open-mixtral-8x7b: We used the following command from the mistralai Python package:
@@
  • Models published in 2024: mistral-small-2402, mistral-large-2402, open-mixtral-8x22b: We used the following command from the mistralai Python package (a JSON mode is available for those model but not for the models published in 2023):
@@
Appendix B.3 Running models with the llama-cpp-python Python library
@@
Appendix B.4 Running models on an Apple computer with the mlx-lm Python library
@@
Appendix Appendix C Messages submitted to LLMs
@@
Appendix C.1 British party manifestos - scaling without explanations about the policy dimensions.
Appendix C.1.1 pre_text
You will be provided with a text from a party manifesto.
Appendix C.1.2 post_text
@@
    Where does this text stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. To compute the score, start by first identifying the parts of the text that are about economic policy. Then, compute the score based on these parts of text. If the text does not refer to economic policy, return "NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
@@
    Where does this text stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. To compute the score, start by first identifying the parts of the text that are about social policy. Then, compute the score based on these parts of text. If the text does not refer to social policy, return "NA"". You will only respond with a JSON object with the key Score. Do not provide explanations
    Immigration policy:
    Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.2 British party manifestos - scaling with explanations about the policy dimensions.
Appendix C.2.1 pre_text
You will be provided with a text from a party manifesto.
Appendix C.2.2 post_text
@@
    Where does this text stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? “Economic” policies deal with all aspects of the economy, including: Taxation; Government spending; Services provided by the government or other public bodies; Pensions, unemployment and welfare benefits, and other state benefits; Property, investment and share ownership, public or private; Interest rates and exchange rates; Regulation of economic activity, public or private; Relations between employers, workers and trade unions. “Left” economic policies tend to favor one or more of the following: High levels of services provided by the government and state benefits, even if this implies high levels of taxation; Public investment. Public ownership or control of sections of business and industry; Public regulation of private business and economic activity; Support for workers/trade unions relative to employers. “Right” economic policies tend to favor one or more of the following: Low levels of taxation, even if this implies low levels of levels of services provided by the government and state benefits; Private investment. Minimal public ownership or control of business and industry; Minimal public regulation of private business and economic activity; Support for employers relative to trade unions/workers. Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. To compute the score, start by first identifying the parts of the text that are about economic policy. Then, compute the score based on these parts of text. If the text does not refer to economic policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
@@
    Where does this text stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? “Social” policies deal with aspects of social and moral life, relationships between social groups, and matters of national and social identity, including: Policing, crime, punishment and rehabilitation of offenders; Immigration, relations between social groups, discrimination and multiculturalism; The role of the state in regulating the social and moral behavior of individuals. “Liberal” social policies tend to favor one or more of the following: Recording the contents of political text on economic and social scales; Policies emphasizing prevention of crime, rehabilitation of convicted criminals; The right of individuals to make personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies penalizing discrimination against particular social groups and/or favoring a multicultural society. “Conservative” social policies tend to favor one or more of the following: Policies emphasizing more aggressive policing, increasing police numbers, conviction and punishment of criminals, building more prisons; The right of society to regulate personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies favoring restriction of immigration, and/or opposing explicit provision of state services for minority cultures. Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. To compute the score, start by first identifying the parts of the text that are about social policy. Then, compute the score based on these parts of text. If the text does not refer to social policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations
    Immigration policy:
    Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? “Immigration” policy relates to all government policies, laws, regulations, and practices that deal with the free travel of foreign persons across the country’s borders, especially those that intend to live, work, or seek legal protection (asylum) in that country. Examples of specific policies that pertain to immigration include the regulation of: work permits for foreign nationals; residency permits for foreign nationals; asylum seekers and their treatment; requirements for acquiring citizenship; illegal immigrants and migrant workers (and their families) living or working illegally in the country. It also includes favorable or unfavorable general statements about immigrants or immigration policy, such as statements indicating that immigration has been good for a country, or that immigrants have forced local people out of jobs, etc. Pro-immigration policies Examples of “pro” immigration positions include: Positive statements about the benefits of immigration, such as economic or cultural benefits; Statements about the moral obligation to welcome asylum seekers; Policies that would improve conditions for asylum seekers and their families; Urging an increase the number of work permits for foreign nationals; Making it possible for illegal immigrants to obtain a legal status or even citizenship; Reducing barriers to immigration generally. Anti-immigration policies Examples of “anti” immigration positions include: Negative statements about consequences of immigration, such as job losses, increased crime, or destruction of national culture; Arguments about asylum seekers abusing the system; Policies to deport asylum seekers and their families; Urging restrictions on the number of work permits for foreign nationals, including points systems; Deporting illegal immigrants and their families; Increasing barriers to immigration generally. Neutral immigration policy statements Examples of neutral statements about immigration policy: Advocating a balanced approach to the problem; Statements about administrative capacity for handling immigration or asylum seekers; Statements that do not take a pro- or anti-immigration stance generally, despite making some statement about immigration. Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return “NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.3 EU legislative speeches - scaling with background explanations.
Appendix C.3.1 pre_text
You will be provided with a text from a debate over policy in the European parliament.
Appendix C.3.2 post_text
@@
We used the following post_text that included background information about the context of the legislative debate (especially the nature of the proposal being discussed), the scale dimension and the two ends of the scale designed to be as close as possible to the instructions given to crowdworkers in Benoit et al. (2016) (note that we reverse-coded the results in the analyses, to be consistent with the scale direction in Benoit et al. (2016)):
@@
Where does this text stand on the ‘pro-subsidy’ to ‘anti-subsidy’ scale, in terms of state subsidies for uncompetitive coal mines? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-subsidy’ and 100 means ‘Extremely anti-subsidy’ ## Background. The debates are taken from a debate in the European Parliament over the ending of state support for uncompetitive coal mines. In general, state aid for national industry in the European Union is not allowed, but exceptions are made for some sectors such as agriculture and energy. At stake here were not only important policy issues as to whether state intervention is preferable to the free market, but also the specific issue for some regions (e.g. Ruhrgebiet in Germany, the north-west of Spain, the Jiu Valley in Romania) where the social and economic impacts of closure would be significant, possibly putting up 100,000 jobs at risk when related industries are considered. Specifically, the debate concerned a proposal by the European Commission to phase out all state support by 2014. Legislation passed in 2002 that allowed for state subsidies to keep non- profitable coal mines running was due to end in 2010. The Commission proposed to let the subsidies end, but to allow limited state support until 2014 in order to soften the effects of the phase-out. A counter proposal was introduced to extend this support until 2018, although many speakers took the opportunity to express very general positions on the issue of state subsidies and energy policy. ## The ‘pro-subsidy’ to ‘anti-subsidy’ scale. The “anti-subsidy" position was the essence of the Commission proposal, which would have let subsidies end while offering limited state aid until 2014 only. Examples of “anti-subsidy" positions: Statements declaring support for the commission position. Statements against state aid generally, for reasons that they distort the market. Arguments in favor of the Commission phase-out date of 2014, rather than 2018. In the strong form, the “pro-subsidy" position, involved rejecting the Commission proposal and favoring continuing subsidies indefinitely. In a weaker form, this involved supporting the compromise to extend limited state support until 2018. Examples of “pro-subsidy" positions: Arguments that keeping the coal mines open to provides energy security. Arguments that coal mines should be kept open to provide employment and other local economic benefits. Preferences for European coal over imported coal, for environmental or safety reasons. ## Instructions. To compute the score, start by first identifying the parts of the text that are about subsidy policy. Then, compute the score based on these parts of text. If the text does not refer to subsidy policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.4 EU legislative speeches - scaling without background explanations.
Appendix C.4.1 pre_text
You will be provided with a text from a debate over policy in the European parliament.
Appendix C.4.2 post_text
@@
Where does this text stand on the ‘pro-subsidy’ to ‘anti-subsidy’ scale, in terms of state subsidies for uncompetitive coal mines? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-subsidy’ and 100 means ‘Extremely anti-subsidy’. To compute the score, start by first identifying the parts of the text that are about subsidy policy. Then, compute the score based on these parts of text. If the text does not refer to subsidy policy, return “NA"". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.5 Tweets
Appendix C.5.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ’right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set score to ‘NA’. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.5.2 post_text
Left empty.
Appendix C.6 Tweets published by each senator
Appendix C.6.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ’Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to“NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix C.6.2 post_text
Left empty.
Appendix C.7 Sets of tweets published by each senators
Appendix C.7.1 pre_text
Appendix C.7.2 post_text
Appendix Appendix D Other Scaling Methods
Appendix D.1 British party manifestos - fine-tuned BERT classifier
@@
We implemented our BERT model using the bert-base-uncased, the TensorFlow machine-learning library and its higher-level wrapper Keras. We used the Adam optimizer to minimize the cross-entropy loss function to fine-tuning the model. This amounts to trying to find the parameters that maximize the likelihood of the true categories in the training data. We used Google Colab with a TPU-equipped virtual machine to train the model. We used the following (standard) parameter values:
  • Batch size: 64
@@
The manifestos used for scaling on the Immigration policy dimension were a separate set of party manifestos. For these, the crowdworkers coded each sentence on a 3 point dimension. We adapted the method just described (in the context of Economic and Social policy dimensions) to this context. This dataset of scaled sentences was much smaller that that for Economic and Social policy, giving a training set of about 11K observation, of validation set of about 1.2K observations and a prediction set of about 12K observations.
Appendix D.2 Tweets - fine-tuned BERT classifier
Note that for tweets, we did not train the models on data produced by human coders, but instead trained the classifiers to predict the party of the Congress member who published the tweet (Democratic Party or Republican Party), following the approach proposed by Le Mens et al. (2023b) and implemented on these Twitter data in Le Mens et al. (2023a). We used the approach used in this latter paper by slightly adapting the jupyter python notebooks made available at on the companion OSF folder (https://osf.io/ta273/).
For the tweets published by the senators of the 117th US congress, we used as a training set the tweets published by senators of the 116th and 117th Congress sessions, ensuring that none of the tweets in the prediction set (100 tweets per senator) was in the training or validation set. We used the categorization probabilities in the two parties to compute the typicality of each tweet in each party (following Le Mens et al. (2023b)) and took the difference in typicalities in the Republican and Democratic parties as the position of the tweet on the left-right ideological spectrum. The position estimate of a senator is the average position of the 100 tweets in the prediction set.

### v3

@@ Appendix A Data
A.2 Senators of the 117th US Congress
A.3 British party manifestos
The texts of the party manifestos used for scaling on the economic and social dimension were obtained from the file CSTA-APSR-master/DataCreated/master.sentences.csv which consisted of a table of sentences.
CFjobsresultsImmigration/f354277_immigration1.csv.
A.4 EU legislative speeches
Appendix B Constructing and submitting prompts to LLMs
@@ B.1 OpenAI API
  • GPT-4o (gpt-4o-2024-05-13), GPT-4 Turbo (gpt-4-turbo-2024-04-09), GPT-4o mini (gpt-4o-mini-2024-07-18) and GPT 3.5 Turbo (gpt-3.5-turbo-0125): We used the following command (a JSON mode is available for these models but not for the earlier models):
@@ B.2 Mistral AI API
  • Models published in 2023: open-mixtral-8x7b: We used the following command from the mistralai Python package:
@@
  • Models published in 2024: mistral-large-2402, mistral-large-2407, open-mixtral-8x22b, open-mistral-nemo-2407: We used the following command from the mistralai Python package (a JSON mode is available for those model but not for the models published in 2023):
@@
B.3 Running models with the llama-cpp-python Python library
@@
B.4 Running models on an Apple computer with the mlx-lm Python library
@@
B.5 Running models on an Huggingface inference endpoint
The Huggingface model sharing platform provides a service called ‘Inference endpoints’ that allows users to run LLMs on dedicated cloud machines provided by Amazon Web Services (AWS), Microsoft Azure or Google Cloud Platform, a location for the machine (e.g., in the US or in Europe). It offers a graphical interface that allows the user to choose an LLM (e.g., Llama 3 70B Instruct), a GPU infrastructure on which to run it (with different amounts of video memory), and various configuration parameters (such as the maximum number of tokens). Once the setup is completed, the user is provided with a custom URL that is then used to send the prompts to the LLM (stored in the variable API_URL in the code snippet below). The cost is per hour (between 1 USD for a machine with 24GB of video memory that can run models such as Llama 3 8B to 16 USD for a machine with 320GB of video memory that can run models such as Llama 3 70B).
headers = {
ΨΨ"Accept" : "application/json",
ΨΨ"Authorization": f"Bearer {HF_TOKEN}",
ΨΨ"Content-Type": "application/json"
Ψ}
def hf_inference_endpoint_generate(payload):
Ψresponse = requests.post(API_URL, headers=headers, json=payload)
ΨΨreturn response.json()
prompt = tokenizer.apply_chat_template(
    messages, tokenize=False, add_generation_prompt=True
response_raw = hf_inference_endpoint_generate({
ΨΨΨΨΨ"inputs": prompt,
ΨΨΨΨΨ"parameters": {
ΨΨΨΨΨΨ"top_p": 0.999,
ΨΨΨΨΨΨ"temperature": 0.001,
ΨΨΨΨΨΨ"max_new_tokens": max_tokens,
ΨΨΨΨΨΨ"do_sample": do_sample,
ΨΨΨΨΨΨ"return_text": False,
ΨΨΨΨΨΨ"return_full_text": False,
ΨΨΨΨΨΨ"return_tensors": False,
ΨΨΨΨΨΨ"clean_up_tokenization_spaces": True
ΨΨΨΨΨ}
ΨΨΨΨ})
Here, HF_TOKEN is a unique identification token users of Huggingface can create in their account settings. The funcion hf_inference_endpoint_generate sends the prompt prompt to the dedicated cloud machine.
Appendix C Messages submitted to LLMs
@@
C.1 Tweets
C.1.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set score to ‘NA’. You will only respond with a JSON object with the key Score. Do not provide explanations.
C.1.2 post_text
Left empty.
C.2 Tweets published by each senator
C.2.1 pre_text
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
C.2.2 post_text
Left empty.
C.3 Sets of tweets published by each senators
C.3.1 pre_text
C.3.2 post_text
C.4 British party manifestos - scaling without explanations about the policy dimensions.
C.4.1 pre_text
You will be provided with a sentence from a party manifesto.
C.4.2 post_text
@@ C.3 Sets of tweets published by each senators
    Where does this sentence stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the sentence does not refer to economic policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
@@
    Where does this sentence stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. If the sentence does not refer to social policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
C.5 British party manifestos - scaling with explanations about the policy dimensions.
C.5.1 pre_text
You will be provided with a sentence from a party manifesto.
C.5.2 post_text
@@
    Where does this sentence stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? “Economic” policies deal with all aspects of the economy, including: Taxation; Government spending; Services provided by the government or other public bodies; Pensions, unemployment and welfare benefits, and other state benefits; Property, investment and share ownership, public or private; Interest rates and exchange rates; Regulation of economic activity, public or private; Relations between employers, workers and trade unions. “Left” economic policies tend to favor one or more of the following: High levels of services provided by the government and state benefits, even if this implies high levels of taxation; Public investment. Public ownership or control of sections of business and industry; Public regulation of private business and economic activity; Support for workers/trade unions relative to employers. “Right” economic policies tend to favor one or more of the following: Low levels of taxation, even if this implies low levels of levels of services provided by the government and state benefits; Private investment. Minimal public ownership or control of business and industry; Minimal public regulation of private business and economic activity; Support for employers relative to trade unions/workers. Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the sentence does not refer to economic policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
@@
    Where does this sentence stand on the ‘liberal’ to ‘conservative’ scale, in terms of social policy? “Social” policies deal with aspects of social and moral life, relationships between social groups, and matters of national and social identity, including: Policing, crime, punishment and rehabilitation of offenders; Immigration, relations between social groups, discrimination and multiculturalism; The role of the state in regulating the social and moral behavior of individuals. “Liberal” social policies tend to favor one or more of the following: Recording the contents of political text on economic and social scales; Policies emphasizing prevention of crime, rehabilitation of convicted criminals; The right of individuals to make personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies penalizing discrimination against particular social groups and/or favoring a multicultural society. “Conservative” social policies tend to favor one or more of the following: Policies emphasizing more aggressive policing, increasing police numbers, conviction and punishment of criminals, building more prisons; The right of society to regulate personal moral choices on matters such as abortion, gay rights, and euthanasia; Policies favoring restriction of immigration, and/or opposing explicit provision of state services for minority cultures. Provide your response as a score between 0 and 100 where 0 means ‘Extremely liberal’ and 100 means ‘Extremely conservative’. If the sentence does not refer to social policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
C.6 EU legislative speeches - scaling with background explanations.
@@
We used the following pre_text that included background information about the context of the legislative debate (especially the nature of the proposal being discussed), the scale dimension and the two ends of the scale designed to be as close as possible to the instructions given to crowdworkers in Benoit et al. (2016):
C.6.1 pre_text
@@
## Summary This task involves reading a sentence from a debate over policy in the European parliament, and judging whether particular statements were for or against a proposed policy. ## Background. The debates are taken from a debate in the European Parliament over the ending of state support for uncompetitive coal mines. In general, state aid for national industry in the European Union is not allowed, but exceptions are made for some sectors such as agriculture and energy. At stake here were not only important policy issues as to whether state intervention is preferable to the free market, but also the specific issue for some regions (e.g. Ruhrgebiet in Germany, the north-west of Spain, the Jiu Valley in Romania) where the social and economic impacts of closure would be significant, possibly putting up 100,000 jobs at risk when related industries are considered. Specifically, the debate concerned a proposal by the European Commission to phase out all state support by 2014. Legislation passed in 2002 that allowed for state subsidies to keep non- profitable coal mines running was due to end in 2010. The Commission proposed to let the subsidies end, but to allow limited state support until 2014 in order to soften the effects of the phase-out. A counter proposal was introduced to extend this support until 2018, although many speakers took the opportunity to express very general positions on the issue of state subsidies and energy policy. ## Your Coding Job. Your key task is to judge individual sentences from the debate according to which of two contrasting positions they supported: - Supporting the rapid phase-out of subsidies for uncompetitive coal mines. This was the essence of the council proposal, which would have let subsides end while offering limited state aid until 2014 only. - Supporting the continuation of subsidies for uncompetitive coal mines. In the strong form, this involved rejecting the Commission proposal and favoring continuing subsidies indefinitely. In a weaker form, this involved supporting the compromise to extend limited state support until 2018. ## The ‘pro-subsidy’ to ‘anti-subsidy’ scale. The “anti-subsidy" position was the essence of the Commission proposal, which would have let subsidies end while offering limited state aid until 2014 only. Examples of “anti-subsidy" positions: Statements declaring support for the commission position. Statements against state aid generally, for reasons that they distort the market. Arguments in favor of the Commission phase-out date of 2014, rather than 2018. In the strong form, the “pro-subsidy" position, involved rejecting the Commission proposal and favoring continuing subsidies indefinitely. In a weaker form, this involved supporting the compromise to extend limited state support until 2018. Examples of “pro-subsidy" positions: Arguments that keeping the coal mines open to provides energy security. Arguments that coal mines should be kept open to provide employment and other local economic benefits. Preferences for European coal over imported coal, for environmental or safety reasons. ## Sentence:
C.6.2 post_text
@@
## Instructions. Where does this sentence stand on the ‘anti-subsidy’ to ‘pro-subsidy’ scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely anti-subsidy’ and 100 means ‘Extremely pro-subsidy.’ If the sentence does not refer to subsidy policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
Appendix D Other Scaling Methods
D.1 British party manifestos - fine-tuned BERT classifier
@@
We implemented our BERT model using the bert-base-uncased, the TensorFlow machine-learning library and its higher-level wrapper Keras. We used the Adam optimizer to minimize the cross-entropy loss function to fine-tuning the model. This amounts to trying to find the parameters that maximize the likelihood of the true categories in the training data. We used Google Colab with a TPU-equipped virtual machine to train the model.
We used the following (standard) parameter values:
  • Batch size: 256
@@
D.2 Tweets - fine-tuned BERT classifier
Note that for tweets, we did not train the models on data produced by human coders, but instead trained the classifiers to predict the party of the Congress member who published the tweet (Democratic Party or Republican Party), following the approach proposed by Le Mens et al. (2023b) and implemented on these Twitter data in Le Mens et al. (2023a). We used the approach used in this latter paper by slightly adapting the jupyter python notebooks made available at on the companion OSF folder (https://osf.io/ta273/). We used the data from that paper. The training set consisted in close to 1M tweets published by US Congress members during the 116th and 117th Congress sessions.
For the tweets published by the senators of the 117th US congress, we used as a training set the tweets published by senators of the 116th and 117th Congress sessions, ensuring that none of the tweets in the prediction set (100 tweets per senator) was in the training or validation set. This resulted in a training set of about 246K tweets. We used the categorization probabilities in the two parties to compute the typicality of each tweet in each party (following Le Mens et al. (2023b)) and took the difference in typicalities in the Republican and Democratic parties as the position of the tweet on the left-right ideological spectrum. The position estimate of a senator is the average position of the 100 tweets in the prediction set.